OpenStudy (anonymous):

An isosceles triangle in which angles B and C have the same measure is shown. Find the measure of each angle whose degree measure us represented with variables. Angle A: x+8y-1 Angle B: 3y+4 Angle C:7x+5

OpenStudy (anonymous):

angle A y=(1-x)/8 plug y into angle B Since angle B = Angle C (3/8)(1-8)+4 = 7x+5 solve for x, then plug in x into your equation for the angles

OpenStudy (anonymous):

I see, how did you get that y=(1-x)/8 equation? ?.?

OpenStudy (anonymous):

angle A = x+8y-1 x+8y-1 =0 x+8y=1 8y=1-x y=(1/8)*(1-x) or y=(1-x)/8

OpenStudy (anonymous):

many times with geometry problems they will have you solve in terms of a variable from one equation to substitute for another variable in another equation

OpenStudy (anonymous):

please, can you explain and show step by step y=(3/8)(1-8)+4=7x+5 and solve for X . I understand Y=1-x/8 Thank you